# Fourier interpolation in dimensions 3 and 4 and real-variable Kloosterman sums

### v1

@@ 5.2. Case $n=0$
Combining Proposition 5.4 and §5.2, for $r\in{\mathbb{R}}$ and $n\in{\mathbb{Z}}$, we conclude that for $s\in[1,1.001]$, $B_{4,n}(r;y,s)$ and $\widetilde{B}_{4,n}(r;y,s)$ are convergent and have the claimed bound.
By dominated convergence theorem, the factor $e^{-\pi y|n|}$ ensures that the Fourier expansions of $F_{2}(\tau,r;s)$ and $\widetilde{F}_{2}(\tau,r;s)$ for $s\in[1,1.001]$ in Lemma 3.4 are both absolutely convergent. By analytic continuation, $F_{2}(\tau,r;1)$ and $\widetilde{F}_{2}(\tau,r;1)$ has the corresponding Fourier expansion given by $B_{4,n}(r;1,y)$ and $\widetilde{B}_{4,n}(r;1,y)$. The explicit formula in (3.17) and (3.18) follows from (3.6), (5.16) and (5.17). Note that when $n\leq-1$, $B_{4,n}(r;y,s)$ and $\widetilde{B}_{4,n}(r;y,s)$ are convergent to 0 when $s\rightarrow 1^{+}$ because of the Gamma factor.
@@ 8.3. Case $n<0$
Similarly to the proof of Proposition 5.4, the following proposition follows from Lemma 8.1, Proposition 8.2, Proposition 8.3, and (5.13). We omit the proof here.
@@
Recall (8.11). Combining the three equations above, we conclude for $\Sigma_{0}$ that
@@
$\lim_{r\rightarrow 0}\sum_{2|c}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c^{3/2}}=\sum_{2|c}\frac{S(0,n,c,\nu_{\Theta}^{3})}{c^{3/2}}.$
We can evaluate the sums of Kloosterman sums above by recalling the calculations in §7.2. Comparing with the quantity $B_{k}(n)$ defined by Bateman (taking $s=3$) [Bat51, (1.02)], one can find that

### v2

@@ 5.2. Case $n=0$
Combining Proposition 5.3 and §5.2, for $r\in{\mathbb{R}}_{+}$ and $n\in{\mathbb{Z}}$, we conclude that for $s\in[1,1.001]$, $B_{4,n}(r;y,s)$ and $\widetilde{B}_{4,n}(r;y,s)$ are absolute convergent and have the claimed bound.
By dominated convergence theorem, the factor $e^{-\pi y|n|}$ ensures that the Fourier expansions of $F_{2}(\tau,r;s)$ and $\widetilde{F}_{2}(\tau,r;s)$ for $s\in[1,1.001]$ in Lemma 3.4 are both absolutely convergent. By analytic continuation, $F_{2}(\tau,r;1)$ and $\widetilde{F}_{2}(\tau,r;1)$ has the corresponding Fourier expansion given by $B_{4,n}(r;1,y)$ and $\widetilde{B}_{4,n}(r;1,y)$. The explicit formula in (3.17) and (3.18) follows from (3.6), (5.7) and (5.8). Note that when $n\leq-1$, $B_{4,n}(r;y,s)$ and $\widetilde{B}_{4,n}(r;y,s)$ are absolute convergent and converge to 0 when $s\rightarrow 1^{+}$ because of the Gamma factor.
@@ 8.3. Case $n<0$
Similarly to the proof of Proposition 5.3, the following proposition follows from Lemma 8.1, Proposition 8.2, partial summation, Proposition 8.3, and by [DLMF, (10.29.1)]
[…]
. We omit the proof here.
@@
Recall (8.12). Combining the three equations above and noting that there are at most $O(m)$ of $q$’s such that $q\leq\pi m$, we conclude for $\Sigma_{0}$ that
@@
$\lim_{r\rightarrow 0}\sum_{2|c}\frac{K(r^{2},n,c,\nu_{\Theta}^{3})}{c^{3/2}}=\sum_{2|c}\frac{S(0,n,c,\nu_{\Theta}^{3})}{c^{3/2}}$
and similarly for the case $2\nmid\widetilde{c}$.
We can evaluate the sums of Kloosterman sums in (9.12) by recalling the calculations in §7.2. Comparing with the quantity $B_{k}(n)$ defined by Bateman (taking $s=3$) [Bat51, (1.02)], one can find that